# Qumode Tensor Networks for False Vacuum Decay in Quantum Field Theory

### v1

@@ Abstract
False vacuum decay in scalar quantum field theory (QFT) plays a central role in cosmology and particle physics but has remained intractable for classical simulation due to its non-perturbative, highly entangled dynamics. We present a novel method for Hamiltonian simulation of false-vacuum decay in a self-interacting scalar QFT, based on a spatial discretisation into a lattice of bosonic modes. Each mode has a continuous or quasi-continuous local Hilbert space whose dimension is large enough to capture the self-interactions in the Hamiltonian. We refer to these local modes generically as “qumodes”, echoing the ideal case in which the lattice would be realised on a continuous-variable quantum computer. Here we show that the same framework can be implemented on a tensor network, in which case the qumodes are quasi-continuous, corresponding to the local sites of the network having a very high dimensional Hilbert space. Using a time-evolving block decimation algorithm with Trotterised time evolution, we capture the real-time dynamics of the scalar field. In (1+1)-dimensions, we initialise a metastable vacuum and observe the coherent entanglement growth that seeds true-vacuum bubble nucleation. Our approach establishes the qumode network as a scalable framework for non-equilibrium scalar QFT phenomena and paves the way for higher-dimensional studies and continuous-variable quantum computing implementations.
@@ I Introduction
Tunnelling in scalar QFT is a cornerstone in our understanding of the dynamics of phase transitions. These processes have profound implications for cosmology, particle physics, and the evolution of the early universe. Indeed, it is thought that they played a central role in setting the initial conditions of the universe, influencing both the large-scale structure of the cosmos and the distribution of matter. Moreover, at least one such transition, namely the electroweak phase transition, is known to have taken place in the early universe. This event could provide explanations for the observed matter-antimatter asymmetry and could generate observable gravitational wave signatures. Thus, these mechanisms provide a crucial theoretical bridge linking microphysical interactions and cosmological scales, and understanding them is essential for reconstructing the history of the cosmos.
@@
In summary, then, quantum entanglement is integral to tunnelling in scalar QFT because it underpins the coherence of quantum states across spatial regions. This coherence is essential for the collective behaviour that characterises the formation of nucleating bubbles, and yet it also makes simulating tunnelling in scalar QFT very challenging: one must effectively remain in the false vacuum while generating sufficient entanglement to produce a coherent instanton state. There have been several studies which have sought to simulate aspects of the tunnelling process, which have provided valuable insights [5, 6, 7, 8, 9, 10, 11, 12, 13, 14]. However, in large part due to this difficulty, they considered restricted setups, such as Ising spin-chain models, rather than a full QFT framework.
@@
$\mathcal{U}(t)=\left[\prod_{i}e^{-iH_{i}\delta t}\right]^{t/\delta t}\leavevmode\nobreak\ ,$ (2)
@@ II Qumode Lattice QFT Framework
${\mathscr{V}}(\varphi)\leavevmode\nobreak\ =\leavevmode\nobreak\ \frac{\lambda}{4!}(\varphi^{2}-\varphi_{0}^{2})^{2}-\varepsilon\varphi\leavevmode\nobreak\ -\leavevmode\nobreak\ {\mathscr{V}}_{0}\leavevmode\nobreak\ ,$ (3)
@@
${\mathscr{V}}(\varphi_{0})\leavevmode\nobreak\ \approx\leavevmode\nobreak\ -2\varphi_{0}\varepsilon\leavevmode\nobreak\ .$ (4)
@@ II.1 Qumode preparation
$H_{\rm QFT}\leavevmode\nobreak\ =\leavevmode\nobreak\ \frac{a}{2}\sum_{n=1}^{N}\left[{\hat{p}}_{n}(t)^{2}+\left(\frac{\hat{q}_{\,{n+1}}(t)-\hat{q}_{n}(t)}{a}\right)^{2}+2\mathscr{V}(\hat{q}_{n})\right]\leavevmode\nobreak\ ,$ (5)
@@
$H_{n}\leavevmode\nobreak\ =\leavevmode\nobreak\ \frac{a}{2}{\hat{p}}_{n}^{2}+a^{-1}{\hat{q}}_{n}^{2}+a\mathscr{V}(\hat{q}_{n})\leavevmode\nobreak\ .$ (6)
@@
$\widetilde{H}_{n}\leavevmode\nobreak\ =\leavevmode\nobreak\ \frac{1}{2}{\hat{p}}_{n}^{2}+\mathscr{V}(\hat{q}_{n})\leavevmode\nobreak\ .$ (7)
@@
$q_{n}^{j}\leavevmode\nobreak\ =\leavevmode\nobreak\ -L+j\times\delta q$ (8)
@@
$|\psi_{n}\rangle\leavevmode\nobreak\ =\leavevmode\nobreak\ \sum_{i=1}^{d}|q^{i}_{n}\rangle\left\langle q^{i}_{n}\mid\psi_{n}\right\rangle\leavevmode\nobreak\ .$ (9)
@@
$\hat{q}_{n}\leavevmode\nobreak$ $=\leavevmode\nobreak\ \begin{pmatrix}q^{1}_{n}&0&\ldots\\ 0&q^{j}_{n}&\\ \vdots&&q^{d}_{n}\end{pmatrix}\leavevmode\nobreak$
$\leavevmode\nobreak\ \leavevmode\nobreak\ \hat{p}_{n}^{2}\leavevmode\nobreak$ $=\leavevmode\nobreak\ -\frac{\hbar^{2}}{\delta q^{2}}\begin{pmatrix}2&-1&\leavevmode\nobreak\ &\leavevmode\nobreak\ \leavevmode\nobreak\ &\ldots\\ -1&2&-1&\leavevmode\nobreak\ &\ldots\\ \leavevmode\nobreak\ &-1&2&\leavevmode\nobreak\ &\ldots\\ \vdots&\leavevmode\nobreak\ &\leavevmode\nobreak\ &\ddots&-1\\ \vdots&\leavevmode\nobreak\ &\leavevmode\nobreak\ &\leavevmode\nobreak\ \leavevmode\nobreak\ -1&2\end{pmatrix}$ (10)
@@
$\widetilde{\mathcal{U}}_{n}(t)\leavevmode\nobreak\ =\leavevmode\nobreak\ e^{-i\widetilde{H}_{n}t}\leavevmode\nobreak\ ,$ (11)
@@
$\widetilde{\mathcal{U}}_{n}(t)\leavevmode\nobreak\ =\leavevmode\nobreak\ \left[e^{-i\hat{p}_{n}^{2}\delta t}e^{-i\mathscr{V}(\hat{q}_{n})\delta t}\right]^{t/\delta t}+\mathcal{O}\left(\delta t^{2}\right)\leavevmode\nobreak\ .$ (12)
@@
In order to test the evolution we start the qumode in the ground state of the SHO around the false vacuum at $\varphi=\varphi_{\rm fv}=-2$. The evolution of a single qumode from this initial state is shown in Fig. 3, where in the first panel we show the evolution of the wavefunction of the continuous qumode for times up to $t=60$, and in the second panel we show the evolution in the discretised qumode with a $d=50$ dimensional Hilbert space. The Trotterised time-evolution has been performed with a time step of $\delta t=0.05$. As we will ultimately be interested in the local field expectation value, $\langle\varphi(x)\rangle\equiv\langle\hat{q}_{n}\rangle$, this approximation is more than adequate for our purposes. (Of course it can be improved by taking smaller Trotter steps and higher dimensionality of local Hilbert spaces.) More generally we see that the evolution displays the expected tunnelling behaviour. But what we also observe is that the tunnelling does not “complete” because, in what is effectively a (1+0)-dimensional field theory, the excess energy has nowhere to dissipate. As we shall see, once the qumodes are tied together into a lattice, this excess energy will ultimately be spent driving bubble expansion.
@@ II.2 Qumode initialisation by adiabatic evolution
${\mathscr{V}}(\varphi)\leavevmode\nobreak\ =\leavevmode\nobreak\ \frac{\omega^{2}}{2}(\varphi-\varphi_{\rm fv})^{2}+{\mathscr{V}}_{1}(\varphi)$ (13)
@@
$\omega^{2}\leavevmode\nobreak\ =\leavevmode\nobreak\ \frac{\lambda\varphi_{0}^{2}}{3}\leavevmode\nobreak\ ,$ (14)
@@
${\mathscr{V}}_{1}(\varphi)\leavevmode\nobreak\ =\leavevmode\nobreak\ \frac{\lambda}{4!}(\varphi+\varphi_{0})^{3}(\varphi-3\varphi_{0})-\varepsilon\varphi\leavevmode\nobreak\ -\leavevmode\nobreak\ {\mathscr{V}}_{0}\leavevmode\nobreak\ .$ (15)
@@
$\langle q_{n}|\psi_{n}(s=0)\rangle\leavevmode\nobreak\ =\leavevmode\nobreak\ \left(\frac{\omega}{\pi}\right)^{1/4}e^{-\omega(q_{n}+\varphi_{0})^{2}/2}\leavevmode\nobreak\ .$ (16)
@@
${\mathscr{V}}_{s}(\hat{q}_{n})\leavevmode\nobreak\ =\leavevmode\nobreak\ \frac{\omega^{2}}{2}(\hat{q}_{n}-\varphi_{\rm fv})^{2}+s\,({\mathscr{V}}_{1}(\hat{q}_{n})+{\mathscr{V}}_{\rm lift}(\hat{q}_{n}))\leavevmode\nobreak\ ,$ (17)
@@
${\mathscr{V}}_{\rm lift}(\hat{q}_{n}))\leavevmode\nobreak\ =\leavevmode\nobreak\ (2+\delta)\,\varphi_{0}\varepsilon e^{-(\hat{q}-\varphi_{0})^{2}}\leavevmode\nobreak\ ,$ (18)
which lifts the minimum at $\hat{q}=\varphi_{0}$ to $+\delta$. Sending the evolution parameter $s$ to one gradually reintroduces all of the remaining ${\mathscr{V}}_{1}$ terms in the potential around the false vacuum at $\hat{q}=\varphi_{\rm fv}=-\varphi_{0}$, while the lift potential ${\mathscr{V}}_{\rm lift}$ prevents the true minimum of the full potential appearing at $\hat{q}_{n}=\varphi_{\rm tv}=+\varphi_{0}$. According to the adiabatic theorem, the wavefunction will remain in the ground state throughout, provided that $s$ evolves sufficiently slowly. The lift potential ensures that the wavefunction does not simply end up in the ground state of the full potential during this process, which of course is centred around $\hat{q}_{n}=+\varphi_{0}$. A comparison of the three potentials is shown in Fig. 4 for the quite extreme case of $\lambda=1$ and $\varepsilon=0.1$.
Figure 4: Initialising the false vacuum state using adiabatic evolution with $\lambda=1$ and $\varepsilon=0.1$: the solid line is the full potential, the dotted line is the SHO approximation around the false vacuum where we begin the qumode in the corresponding SHO groundstate, and the dashed line is the full potential plus lift potential, where here we have taken $\delta=4$. During the adiabatic evolution the potential is gradually changed from the dotted SHO potential to the dashed potential.
Figure 5: The SHO groundstate versus the true ground state for the full potential, and the “false vacuum ground state”. Both of these states were found by adiabatically evolving from the SHO ground state. As might have been anticipated from the shape of the full potential compared to the SHO potential, the “false vacuum ground state” is shifted to the right and broadened compared to the SHO groundstate.
@@
$\widetilde{\mathcal{W}}_{n}\leavevmode\nobreak\ =\leavevmode\nobreak\ \prod_{r=1}^{M}\left[e^{-i\hat{p}_{n}^{2}\delta t}e^{-i\frac{\omega^{2}}{2}(\hat{q}_{n}-\varphi_{\rm fv})^{2}\delta t}\leavevmode\nobreak\ {\widetilde{\mathcal{V}}}^{\,r}_{n}\right]\leavevmode\nobreak\ ,$ (19)
@@
$\widetilde{\mathcal{V}}_{n}\leavevmode\nobreak\ =\leavevmode\nobreak\ e^{-i{\mathscr{V}}_{1}(\hat{q}_{n})\delta s\,\delta t}\leavevmode\nobreak\ ,$ (20)
and where the Trotter step $\delta t$, which is chosen to be appropriately small, incorporates normal time evolution. Note that this requires only the single non-linear gate $\widetilde{\mathcal{V}}_{n}$ for a qumode, and in principle one need only determine and record the correct initial qumode state once for any given potential (although in practice it is possible to simply include this procedure as part of the qumode initialisation of each QFT).
@@
$t_{\rm tot}\leavevmode\nobreak\ \gg\leavevmode\nobreak\ \frac{1}{(\mathscr{E}_{1}-\mathscr{E}_{0})_{\rm min}^{2}}\leavevmode\nobreak\ ,$ (21)
@@
We show an example of this procedure in Fig. 5 for the quartic potential of Fig. 4 (which is a somewhat extreme case). The curves shown are the initial SHO ground state which is centred around $\varphi=-\varphi_{0}$, the true ground state which is of course centred around $\varphi=\varphi_{0}$, found by adiabatically evolving to the full potential (requiring at least $10^{4}$ Trotter steps in order to remain adiabatic in this case), and the “false vacuum ground state” found by adiabatically evolving to the full $+$ lift potential.

### v2

@@ Abstract
False vacuum decay in scalar quantum field theory (QFT) is a cornerstone of early Universe cosmology and high energy physics, yet its real-time dynamics is essentially inaccessible to classical computation due to its non-perturbative, highly entangled dynamics. We introduce a general Hamiltonian framework for simulating full interacting QFTs, using a spatial lattice of continuos-variable “qumodes” – bosonic local oscillators whose high-dimensional local Hilbert space faithfully captures interacting field dynamics. This construction is rooted in continuous-variable quantum computing (CVQC), and provides a unified platform spanning efficient classical tensor-network methods and emerging photonic quantum hardware. The first key advance of this work is a robust and scaleable method for preparing the QFT in its correct initial vacuum state. We develop an imaginary-time preparation algorithm tailored to qumode lattices, that efficiently projects onto the vacuum even in strongly coupled regimes. This provides a controllable starting point for studying nonperturbative dynamics such as tunnelling and real-time decay. Building on this, we use a time-evolving block decimation algorithm to capture the real-time dynamics of the scalar field. Our second key advance is the identification and excitation of the negative fluctuation mode of the bounce configuration on the qumode lattice. A small displacement along this mode produces the expected tachyonic growth, driving fully coherent bubble nucleation without requiring classically supercritical seeds. This demonstrates that the qumode lattice captures non-perturbative quantum dynamics that lie beyond the classical treatments. Our results establish the qumode network as a scalable framework for non-equilibrium scalar QFT phenomena and pave the way for higher-dimensional studies and continuous-variable quantum computing implementations.
@@ I Introduction
The purpose of this paper is to establish the first framework that enables direct, real-time simulation of non-perturbative processes, within a full scalar quantum field theory (QFT). By formulating the problem in the Hamiltonian picture and representing the field as a lattice of continuous-variable modes, our approach captures non-perturbative dynamics beyond the reach of semiclassical and path-integral techniques. Furthermore, by exploiting continuous-variable quantum computing (CVQC) techniques, this approach provides an efficient route towards quantum simulation of scalar QFTs, reducing the required quantum volume in comparison to discrete-variable quantum computing (DVQC) platforms. In this paper, we focus on simulating the real-time dynamics of false-vacuum decay, establishing a new route to studying phase transitions from first principles, bridging fundamental quantum field dynamics and quantum simulation platforms.
False-vacuum decay in scalar QFT is a cornerstone in our understanding of the dynamics of phase transitions. These processes have profound implications for cosmology, particle physics, and the evolution of the early universe. Indeed, it is thought that they played a central role in setting the initial conditions of the universe, influencing both the large-scale structure of the cosmos and the distribution of matter. Moreover, at least one such transition, namely the electroweak phase transition, is known to have taken place in the early universe. This event could provide explanations for the observed matter-antimatter asymmetry and could generate observable gravitational wave signatures. Thus, these mechanisms provide a crucial theoretical bridge linking microphysical interactions and cosmological scales, and understanding them is essential for reconstructing the history of the cosmos.
@@
In summary, then, quantum entanglement is integral to tunnelling in scalar QFT because it underpins the coherence of quantum states across spatial regions. This coherence is essential for the collective behaviour that characterises the formation of nucleating bubbles, and yet it also makes simulating tunnelling in scalar QFT very challenging: one must effectively remain in the false vacuum while generating sufficient entanglement to produce a coherent instanton state. There have been several studies which have sought to simulate aspects of the tunnelling process, which have provided valuable insights [5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15]. However, in large part due to this difficulty, they considered restricted setups, such as Ising spin-chain models, rather than a full QFT framework.
@@
$\mathcal{U}(t)=\left[\prod_{i}e^{-iH_{i}\delta t}\right]^{t/\delta t}\penalty 10000\ ,$ (2)
@@
A crucial advance we make here is an algorithm for initial state preparation of the full QFT. This is done by running the same TEBD through imaginary time, which projects onto the lowest energy state, arranged to be the metastable vacuum state.
@@ II Qumode Lattice QFT Framework
${\mathscr{V}}(\varphi)\penalty 10000\ =\penalty 10000\ \frac{\lambda}{4!}(\varphi^{2}-\varphi_{0}^{2})^{2}-\varepsilon\varphi\penalty 10000\ -\penalty 10000\ {\mathscr{V}}_{0}\penalty 10000\ ,$ (3)
@@
${\mathscr{V}}(\varphi_{0})\penalty 10000\ \approx\penalty 10000\ -2\varphi_{0}\varepsilon\penalty 10000\ .$ (4)
@@ II.1 Qumode preparation
$H_{\rm QFT}\penalty 10000\ =\penalty 10000\ \frac{a}{2}\sum_{n=1}^{N}\left[{\hat{p}}_{n}(t)^{2}+\left(\frac{\hat{q}_{\,{n+1}}(t)-\hat{q}_{n}(t)}{a}\right)^{2}+2\mathscr{V}(\hat{q}_{n})\right]\penalty 10000\ ,$ (5)
@@
$H_{n}\penalty 10000\ =\penalty 10000\ \frac{a}{2}{\hat{p}}_{n}^{2}+a^{-1}{\hat{q}}_{n}^{2}+a\mathscr{V}(\hat{q}_{n})\penalty 10000\ .$ (6)
@@
$\widetilde{H}_{n}\penalty 10000\ =\penalty 10000\ \frac{1}{2}{\hat{p}}_{n}^{2}+\mathscr{V}(\hat{q}_{n})\penalty 10000\ .$ (7)
@@
$q_{n}^{j}\penalty 10000\ =\penalty 10000\ -L+j\times\delta q$ (8)
@@
$|\psi_{n}\rangle\penalty 10000\ =\penalty 10000\ \sum_{i=1}^{d}|q^{i}_{n}\rangle\left\langle q^{i}_{n}\mid\psi_{n}\right\rangle\penalty 10000\ .$ (9)
@@
$\hat{q}_{n}\penalty 10000\$$=\penalty 10000\ \begin{pmatrix}q^{1}_{n}&0&\ldots\\ 0&q^{j}_{n}&\\ \vdots&&q^{d}_{n}\end{pmatrix}\penalty 10000\$
$\penalty 10000\ \penalty 10000\ \hat{p}_{n}^{2}\penalty 10000\$$=\penalty 10000\ -\frac{\hbar^{2}}{\delta q^{2}}\begin{pmatrix}2&-1&\penalty 10000\ &\penalty 10000\ \penalty 10000\ &\ldots\\ -1&2&-1&\penalty 10000\ &\ldots\\ \penalty 10000\ &-1&2&\penalty 10000\ &\ldots\\ \vdots&\penalty 10000\ &\penalty 10000\ &\ddots&-1\\ \vdots&\penalty 10000\ &\penalty 10000\ &\penalty 10000\ \penalty 10000\ -1&2\end{pmatrix}$ (10)
@@
$\widetilde{\mathcal{U}}_{n}(t)\penalty 10000\ =\penalty 10000\ e^{-i\widetilde{H}_{n}t}\penalty 10000\ ,$ (11)
@@
$\widetilde{\mathcal{U}}_{n}(t)\penalty 10000\ =\penalty 10000\ \left[e^{-i\hat{p}_{n}^{2}\delta t}e^{-i\mathscr{V}(\hat{q}_{n})\delta t}\right]^{t/\delta t}+\mathcal{O}\left(\delta t^{2}\right)\penalty 10000\ .$ (12)
@@
In order to test the evolution we start the qumode in the ground state of the SHO with the mass corresponding to $\omega$, which is a Gaussian around the false vacuum at $\varphi=\varphi_{\rm fv}=-2$. The evolution of a single qumode from this initial state is shown in Fig. 3, where in the first panel we show the evolution of the wavefunction of the continuous qumode for times up to $t=60$, simulated using qibo [62], and in the second panel we show the evolution in the discretised bosonic mode with a $d=50$ dimensional Hilbert space. The Trotterised time-evolution has been performed with a time step of $\delta t=0.05$. In the lower panel, minor deviations due to the finite truncation of the local Hilbert space are visible as a slight smoothing of the probability density at later times relative to the exact case. Despite this, it is evident that the method is resilient to Trotter error even for long evolution times. Therefore, we expect the full QFT simulations to exhibit similar robustness, with Trotter errors remaining small for sufficiently small $\delta t$.
As we will ultimately be interested in the local field expectation value, $\langle\varphi(x)\rangle\equiv\langle\hat{q}_{n}\rangle$, this approximation is more than adequate for our purposes. (Of course it can be improved by taking smaller Trotter steps and higher dimensionality of local Hilbert spaces.) More generally we see that the evolution displays the expected tunnelling behaviour. But what we also observe is that the tunnelling does not “complete” because, in what is effectively a (1+0)-dimensional field theory, the excess energy has nowhere to dissipate. As we shall see, once the qumodes are tied together into a lattice, this excess energy will ultimately be spent driving bubble expansion.
@@ II.2 Qumode initialisation by adiabatic evolution
${\mathscr{V}}(\varphi)\penalty 10000\ =\penalty 10000\ \frac{\omega^{2}}{2}(\varphi-\varphi_{\rm fv})^{2}+{\mathscr{V}}_{1}(\varphi)$ (13)
@@
$\omega^{2}\penalty 10000\ =\penalty 10000\ \frac{\lambda\varphi_{0}^{2}}{3}\penalty 10000\ ,$ (14)
@@
${\mathscr{V}}_{1}(\varphi)\penalty 10000\ =\penalty 10000\ \frac{\lambda}{4!}(\varphi+\varphi_{0})^{3}(\varphi-3\varphi_{0})-\varepsilon\varphi\penalty 10000\ -\penalty 10000\ {\mathscr{V}}_{0}\penalty 10000\ .$ (15)
@@
$\langle q_{n}|\psi_{n}(s=0)\rangle\penalty 10000\ =\penalty 10000\ \left(\frac{\omega}{\pi}\right)^{1/4}e^{-\omega(q_{n}+\varphi_{0})^{2}/2}\penalty 10000\ .$ (16)
@@
${\mathscr{V}}_{s}(\hat{q}_{n})\penalty 10000\ =\penalty 10000\ \frac{\omega^{2}}{2}(\hat{q}_{n}-\varphi_{\rm fv})^{2}+s\,({\mathscr{V}}_{1}(\hat{q}_{n})+{\mathscr{V}}_{\rm lift}(\hat{q}_{n}))\penalty 10000\ ,$ (17)
@@
${\mathscr{V}}_{\rm lift}(\hat{q}_{n}))\penalty 10000\ =\penalty 10000\ (2+\delta)\,\varphi_{0}\varepsilon e^{-(\hat{q}-\varphi_{0})^{2}}\penalty 10000\ ,$ (18)
which lifts the minimum at $\hat{q}=\varphi_{0}$ to $\delta$. Sending the evolution parameter $s$ to one gradually reintroduces all of the remaining ${\mathscr{V}}_{1}$ terms in the potential around the false vacuum at $\hat{q}=\varphi_{\rm fv}=-\varphi_{0}$, while the lift potential ${\mathscr{V}}_{\rm lift}$ prevents the true minimum of the full potential appearing at $\hat{q}_{n}=\varphi_{\rm tv}=+\varphi_{0}$. According to the adiabatic theorem, the wavefunction will remain in the ground state throughout, provided that $s$ evolves sufficiently slowly. The lift potential ensures that the wavefunction does not simply end up in the ground state of the full potential during this process, which of course is centred around $\hat{q}_{n}=+\varphi_{0}$. A comparison of the three potentials is shown in Fig. 4 for the quite extreme case of $\lambda=1$ and $\varepsilon=0.1$.
Figure 4: Initialising the false vacuum state with $\lambda=1$ and $\varepsilon=0.1$: the solid line is the full potential, the dotted line is the SHO approximation around the false vacuum where we begin the qumode in the corresponding SHO groundstate, and the dashed line is the full potential plus lift potential, where here we have taken $\delta=4$. During the adiabatic preparation of the vacuum, the potential is gradually changed from the dotted SHO potential to the dashed potential. For imaginary-time TEBD preparation of the false vacuum, the potential is set as the dashed potential.
Figure 5: The SHO groundstate versus the true ground state for the full potential, and the “false vacuum ground state”. Both of these states were found by adiabatically evolving from the SHO ground state. As might have been anticipated from the shape of the full potential compared to the SHO potential, the “false vacuum ground state” is shifted to the right and broadened compared to the SHO ground state. The simulation has been run for $t=300$ with a $\delta t=0.01$.
@@
$\widetilde{\mathcal{W}}_{n}\penalty 10000\ =\penalty 10000\ \prod_{r=1}^{M}\left[e^{-i\hat{p}_{n}^{2}\delta t}e^{-i\frac{\omega^{2}}{2}(\hat{q}_{n}-\varphi_{\rm fv})^{2}\delta t}\penalty 10000\ {\widetilde{\mathcal{V}}}^{\,r}_{n}\right]\penalty 10000\ ,$ (19)
@@
$\widetilde{\mathcal{V}}_{n}\penalty 10000\ =\penalty 10000\ e^{-i{\mathscr{V}}_{1}(\hat{q}_{n})\delta s\,\delta t}\penalty 10000\ ,$ (20)
and where the Trotter step $\delta t$, which is chosen to be appropriately small, incorporates normal time evolution. Note that this requires only the single non-linear gate $\widetilde{\mathcal{V}}_{n}$ for a qumode, which provides the ramping, and in principle one need only determine and record the correct initial qumode state once for any given potential (although in practice it is possible to simply include this procedure as part of the qumode initialisation of each QFT).
@@
$t_{\rm tot}\penalty 10000\ \gg\penalty 10000\ \frac{1}{(\mathscr{E}_{1}-\mathscr{E}_{0})_{\rm min}^{2}}\penalty 10000\ ,$ (21)
where $\mathscr{E}_{0}(t)$ and $\mathscr{E}_{1}(t)$ are the eigenenergies of the ground state and first excited state at time $t$. This implies that the ground state of the full potential is more difficult to find using this method find than the “ground state of the false vacuum”. Indeed for the former, it is reasonable to suppose that ${\mathscr{E}_{1}-\mathscr{E}_{0}}\,\approx\,4\varepsilon$, which is parametrically small: thus we require
$M\penalty 10000\ \gg\penalty 10000\ 1/\varepsilon^{2}\delta t\penalty 10000\ ,$ (22)
@@
We show an example of this procedure in Fig. 5 for the quartic potential of Fig. 4 (which is a somewhat extreme case). The curves shown are the initial SHO ground state which is centred around $\varphi=-\varphi_{0}$, the true ground state which is of course centred around $\varphi=\varphi_{0}$, found by adiabatically evolving to the full potential (requiring at least $10^{4}$ Trotter steps in order to remain adiabatic in this case), and the “false vacuum ground state” found by adiabatically evolving to the full ++ lift potential.